OpenStudy (anonymous):

what are all real solutions to (square root of x+5)+(square root of x-3)=9

sam (.sam.):

\[(\sqrt{x}+5)+(\sqrt{x}-3)=9~~~?\]

OpenStudy (anonymous):

i think he means the 5 and 3 are under the root.

sam (.sam.):

\[Add\] \[2\sqrt{x}+2=9\] \[2\sqrt{x}=7\] \[\sqrt{x}=\frac{7}{2}\] Square both sides, \[x=\frac{49}{4}\]

sam (.sam.):

If square root of( x+5)+square root of( x-3)=9 \[\sqrt{x+5}+\sqrt{x-3}=9\] \[(\sqrt{x+5}+\sqrt{x-3})^2=9^2\] \[(\sqrt{x+5}+\sqrt{x-3})(\sqrt{x+5}+\sqrt{x-3})=81\] Expand, \[(x+5)+2\sqrt{x+5}\sqrt{x-3}+(x-3)=81\] \[2x+2+2\sqrt{(x+5)(x-3)}=81\] \[2x+2\sqrt{(x+5)(x-3)}=79\] expand , then square both sides again